OpenStudy (anonymous):

2 √(x + 7) – 10 = – 8 2 √(x + 7) – 10 = – 8 @Mathematics

OpenStudy (anonymous):

A. x = 74 B. x = –4 C. x = 32 D. x = –6

OpenStudy (anonymous):

2 sqrt(x + 7) - 10 = -8 add 10 to both sides 2 sqrt(x + 7) = 2 divide both sides by 2 sqrt(x + 7) = 1 square both sides x + 7 = 1 subtract 7 from both sides x = -6

OpenStudy (anonymous):

Thank you